answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

How many bit strings of length 10 have exactly three 0s?

Basically, this is the same as finding the number of distinct ways of arranging seven 1s and 3 0s. That is (10!/7!3!) = (10*9*8)/(3*2*1) = 120.

There are 120 bit strings of length 10 with exactly three 0s.

How can you approximate a binomial distribution to a poison distribution when the number of binomial trials became large enough?

The Poisson distribution with parameter np will be a good approximation for the binomial distribution with parameters n and p when n is large and p is small. For more details See related link below

What number will come up the most when rolling dice?

since there are 6 sides to it any of them could show up, 1,2,3,4,5, and 6 could all show up unlikely

What is the number of outcomes if you roll 2 dice 5 times?

The number of outcomes you would have is number of outcomes x number of times done.

Or in this case, number (of outcomes on the dice) x 5 x 5 x 5 x 5 x 5 because you rolled the dice 5 times

What is the probability that one of two consecutive cards is a spade?

From a standard deck of playing cards, there are 52 cards. There are 13 cards for each Spades, Hearts, Clubs and Diamonds. For the first card, the probability of getting a Spade is 13/52.

If you're going to replace the first card into the deck, then the probability of second card being a Spade is also 13/52

Therefore, 13/52 x 13/52 = 1/16 (one in every sixteen tries)

If you're not going to replace the first card into the deck, the remaining deck will be left with a total of 51 cards and 12 Spades. Therefore the probability of second card being a Spade is 12/51

Lastly, 13/52 x 12/51 = 1/17 (one in every seventeen tries)

What is the probability of rolling at least one six with one roll of a pair of dice?

To find the probability of rolling at least one six with one roll of a pair of dice, we need to know two things. First we need to know how many different possible outcomes there are when we roll the dice one time. We know that there are 36 different possible outcomes. They are 1-1, 1-2, 1-3,..., 6-4, 6-5 and 6-6.

We also need to know how many different times at least one six can appear. These will include, 1-6, 2-6, 3-6, 4-6, 5-6, 6-6, 6-1, 6-2, 6-3, 6-4, and 6-5, for a total of 11 possible outcomes with at least one six. (Note that the 6-6 possible outcome is included in the list because it fits the requirement of "at least one six" with a given single pass of the dice.

The odds of rolling at least one six in one pass with a pair of dice are 11 in 36. To find the probability of this, we need to divide the 11 by the 36, and this gives us a probability of 11/36 or 0.3055555....

Remember that probability is a ratio, and it's the ratio of the number of possible times a given outcome can appear to the total number of possible outcomes that could appear. This ratio will range from zero to one, inclusive. A probability of zero means that a given outcome cannot occur. Like rolling a 13 with one pass of a regular pair of dice. That can't happen, so the probability of it happening is zero.

We also understand that if the probability of something happening is one, then that outcome must happen. Like the probability of rolling an odd or an even number. We know that an odd or an even number must appear, so the probability of that happening is one.

What is the probability of flipping a coin twice and getting two heads?

The probability is 25%.

The probability of flipping a coin once and getting heads is 50%. In your example, you get heads twice -- over the course of 2 flips. So there are two 50% probabilities that you need to combine to get the probability for getting two heads in two flips.

So turn 50% into a decimal --> 0.5

Multiply the two 50% probabilities together --> 0.5 x 0.5 = 0.25.

Therefore, 0.25 or 25% is the probability of flipping a coin twice and getting heads both times.

Why is theoretical probability called theoretical probability?

It is the theory of what might happen, but not actually what happens. In theory, if you spin a coin 100 times, it should come up on heads 50 times, as there is a 1 in 2 chance of you getting heads on each spin. If you actually do spin a coin 100 times, the total of heads is the experimental probability, so what you actually get. That may not be 50. It is likely to be close to 50 though.

What is the theoretical probability of rolling a 5 on a number cube if it is rolled 300 times?

The theoretical probability of rolling a 5 on a standard six sided die is one in six. It does not matter how many times you roll it, however, if you roll it 300 times, the theoretical probability is that you would roll a 5 fifty times.

What is the probability of rolling a number equal to or higher than 5 on a number cube?

There is only one other number higher than five on the standard six sided number cube.

That being said, their is only a one in two, or 50% chance that you will roll equal or higher than five. But only if the five and six are what's relative. It you count the numbers one through four, then you only have a two in six, or one in three (33.3%) chance of rolling five or six.

How many ways can you arrange 4 squares?

Since we're arranging squares, I'm assuming the arrangements are all going to be only two-dimensional and no stacking is allowed. If you have played tetris before, you'll know that can be arranged into :

1) A straight line

2) A 2 by 2 square

3) A T-shape

4) A L-shape

5) An inverted L-shape

6) A ladder

7) An inverted ladder

That's all the seven different arrangements for four squares.

What is the probability of getting either a sum of 5 or at least one 5 in the roll of a pair of dice?

Calculating getting a five on a dice:

There are 6 sides to a dice so there would be 12 in a pair of dice. So you would have a 1 in 6 chance of rolling a 5.

Calculating getting a sum of five:

There are a total of 18 combinations you can get when rolling two dices. 1+1, 1+3, 1+4, 1+5, 1+6, 2+2, 2+3, 2+4, 2+5, 2+6, 3+3, 3+5, 3+6, 4+4, 4+5, 5+5, 6+4, and 6+6 (I eliminated ones which were equal to others such as eliminating 6+2 because it is the same as 2+6). Out of the combination of numbers, only two are able to sum up to five (2+3 and 1+4). This equals about an 11.1% chance of rolling two which equal 5 when added together.

How many black kings in a 52 deck of cards?

There are 2 - the king of clubs, and the king of spades.

How many 6 of clubs are there in a deck of cards?

Their are Four 6 0f clubs in a deck of playing cards.

Just like everything else in it, every thing is in fours.